OpenStudy (anonymous):

Which is 10^–3 written in standard form? A.–30 B.0.0001 C.0.001 D.1000

jimthompson5910 (jim_thompson5910):

use a calculator to evaluate 10^(-3) or 1/(10^3) = 1/1000 = ???

OpenStudy (anonymous):

0.001

OpenStudy (anonymous):

Use this website for questions like that it'll show you how it's done also :3 http://www.mathpapa.com/algebra-calculator.html

OpenStudy (anonymous):

Thank you :D

OpenStudy (anonymous):

You're welcome! ^^